### v1

@@ Abstract
Revert protection is a feature provided by some blockchain platforms that prevents users from incurring fees for failed transactions. This paper explores the economic implications and benefits of revert protection, in the context of priority auctions and maximal extractable value (MEV). We develop an equilibrium game theoretic model that captures the behavior of users (MEV searchers) bidding to have their transaction included ahead of others, in an environment where only a single transaction will succeed in realizing the common value of an opportunity, and in settings both with and without revert protection. Our model applies to a broad range of settings, including Layer 1 (L1) blockchains (e.g., Ethereum mainnet) and Layer 2 (L2) blockchains, and auctions such as “bundle auctions” (on L1s) or priority ordering auctions (on L2s).
We establish that, in the absence of revert protection, users will employ randomized strategies to mitigate the impact of paying for failed transactions. This will ultimately result in less auction revenue, despite the fact that failed transactions still pay fees. Our results quantify in closed form how revert protection enhances auction revenue, and also improves market efficiency and provides for more efficient use of blockspace, as a function of the underlying parameters (the value of the MEV opportunity, the base fee, the revert penalties, and the number of participating agents).
@@ 1 Introduction and Background
Revert protection is a feature for blockchains that block builders can provide where they exclude transactions that would otherwise fail, protecting users from paying fees for failed transactions. However, failed transactions would pay fees, which often accrue to the block builder. So is it in the interest of those builders to offer revert protection?
In this paper, we make the case that it is, and quantify by how much. We study the implications of revert protection on blockchain platforms, and particularly on the auctions that many of those platforms use to capture and allocate maximal extractable value (MEV). We show that in many relevant settings, revert protection is beneficial for auctioneer revenue, as well as for other relevant outcomes like price discovery and blockspace efficiency.
The analysis is complicated by the wide variety of rules — across a diverse variety of L2s, as well as in the complex Ethereum block builder market — for how transactions are selected for inclusion and ordered, how fees are collected, and who those fees go to, as well as how applications behave under those rules. We propose a model that can be parameterized to cover a wide variety of these settings, and use that model to solve for the equilibrium behavior of MEV searchers as a function of those parameters.
@@ 1.1 Revert Protection
On Ethereum and similar blockchains, users interact with smart contracts using atomic transactions. During execution of a transaction, one of those smart contracts may trigger a “revert,” which causes the entire transaction to fail. Normally, if the transaction is included, the user who sent the reverting transaction is still charged some transaction fee, even though the transaction has no other effect on the blockchain state.
The fee charged is typically the product of the transaction’s gas price and the gas used by the transaction before it either completes or reverts. The gas price, in turn, can be separated into the “base fee”—which can be thought of as a flat fee that must be paid by any transaction in the block—and the “priority fee”— an additional fee that is paid to the builder and often affects ordering as well as inclusion (particularly when blocks are full).
The gas used in the transaction is a function of how much of the transaction was executed—a transaction that reverts will usually use only a portion of the gas that it would have used if it succeeded. For example, on an automated market maker, typical reverting transactions may use only around 10–-20% of the gas that a succesful transaction would use.
Some block builders implement a feature where they exclude failed transactions entirely, thus saving the user from having to pay a fee for it. This feature is known as “revert protection.” This feature improves user experience, since users only have to pay a fee if their transaction succeeds. However, it also has significant effects on the behavior of the profit-seeking bots known as “searchers,” which we now describe.
1.2 Searchers, Auctions, and Block Building
Public blockchains have abundant opportunities for MEV. For example, a major type of MEV is the arbitraging of prices between decentralized and centralized exchanges, also called “CEX-DEX arbitrage,” as discussed in [2] and [8]. On these blockchains, independent actors known as “searchers” seek out MEV opportunities and compete to fill them.
Some block builders use different types of auctions to allocate these opportunities to searchers. For a given MEV opportunity, each transaction trying to claim it can be thought of as a “bid.” Generally, only one transaction trying to claim a given opportunity will succeed in extracting the value from it. Revert protection is therefore very relevant for these auctions, because it protects bidders on a given MEV opportunity from having to make a payment for failed bids. Two examples of block building auctions include the “bundle auctions” run by builders such as Flashbots on Ethereum mainnet, and “priority auctions” run by sequencers on Ethereum L2s.
@@ 1.2 Searchers, Auctions, and Block Building
On Ethereum mainnet, blocks are typically built by profit-maximizing “builders.” [^1] Block builders also generally play the role of bidders themselves in the “MEV-Boost” auction, in which they bid to have their block included by the current proposer, as discussed in [4]. Since in that auction, “bids” are complete blocks, revert protection is not as relevant for it, and we will mostly set it aside for purposes of this paper. Its implications for the bundle auction are noted in footnotes below. Many of these builders run “bundle auctions” in which they allow any searcher to submit transactions, and use those searchers’ bids as a factor when deciding which transactions to include and how to order them.
@@
There are two ways bids can be expressed in this auction. First, through the priority fee on the transaction. Second, by making a payment directly as part of the logic of the transaction (by calling coinbase.transfer(...)), as discussed in [3]. The relevant difference between these methods of payment is that (some portion of) priority fees are paid even if the transaction reverts, whereas if the transfer is made as part of the transaction, then it will be conditional on whether the transaction succeeds.
@@
On Ethereum mainnet, the different components of gas price — base and priority fees — are paid in ETH and accrue to different users. Base fees are burned (meaning they ultimately accrue to all ETH holders, rather than the builder). Priority fees accrue to the block builder. [^2] Note that these bids — whether paid through the priority fee or through a transfer to the block’s coinbase address, technically go to the block proposer, not the block builder who is assembling the block. However, since they reduce the amount that the builder has to pay to the proposer to win the MEV-Boost auction, we can think of these payments as a value transfer to the builder.
Many builders, such as Flashbots [5], provide revert protection for transactions submitted to them, even though it is not required by the Ethereum protocol and they might receive more in transaction fees by including it. Our results in this paper help explain this choice by showing that it likely increases their expected revenue in equilibrium.
Priority ordering auctions on L2.
@@
On Layer 2 blockchains today, blocks are typically built by a single sequencer, which often follows a deterministic algorithm for transaction inclusion and ordering. One of the most popular algorithms for this is priority ordering, in which transactions are ordered in descending order of their gas price.
@@
Priority ordering can be thought of as an auction in which transactions bid with the discretionary part of their gas price (i.e., their priority fee) to be included earlier in the chain. By default, this means that certain kinds of MEV (such as top-of-block CEX-DEX arbitrage) will generally accrue to the sequencer through priority fees. This means that when such transactions fail, they will still pay some of their priority fee to the sequencer. However, as discussed in [10], there is a technique called MEV taxes that applications can use to capture all but a negligible portion of the value that would otherwise be paid through priority fees. Since MEV taxes are paid as part of the transaction and revert if the transaction fails, these fees will only be paid if the transaction succeeds.
@@
Even just within these two settings of L1 block builders and L2 priority-ordered sequencers, we now need to consider at least seven cases that differ in who fees go to and how much is paid when the transaction reverts. Table 1 shows the differences between these cases.
@@
Table 1: Transaction Fee Distribution and Revert Behavior
Our model is general enough to compute expected base fee and remaining fee (meaning priority fee, transfer to the coinbase address, or MEV tax) for all of these, by setting different parameters.
1.3 Execution Costs of Revert Protection
This paper is primarily concerned with the potential effects of revert protection, not with its implementation. We should note that feasible revert protection is a difficult technical challenge. Knowing whether a transaction will revert usually requires a builder to execute part of that transaction, which consumes some computational resources. If the builder does not charge the sender anything for that reverted transaction, it may be possible to denial-of-service (DOS) attack the builder with an overwhelming number of reverting transactions.
There are some mitigations to this kind of attack. In some settings, the builder may be able to use out-of-band spam prevention techniques (such as IP blocking) to make this DOS attack unfeasible. For certain special cases—including the auction-like use cases considered in this paper—it may also be possible for the builder to determine statically whether a transaction will succeed or fail.
For purposes of this paper, we leave those challenges out of scope, and assume that—at least for the use cases discussed here—the builder is able to implement revert protection with negligible cost.
@@ 1.4 Our Contributions
  • We introduce a novel, unified game theoretic model that can be used to analyze revert protection in a variety of settings such as L1 block builders, L2 priority-ordered sequencers, MEV taxes, etc.
  • We are able to solve for equilibria in our model in closed form in the model parameters: the value of the MEV opportunity, the base fee, the revert penalty parameters, and the number of participating agents.
  • Using our equilibrium model, we can quantify the benefits of revert protection versus not offering revert protection:
    • Revert protection offers higher auction revenue. Both with and without revert protection, in our setting, the auctioneer can extracts all value when the auction clears. However, in the absence of revert protection, agents randomize when they participate, and there is a non-zero probability that the auction does not clear and value is lost. This results in reduced sequencer revenue.
    • In the context of automated market making, revert protection offers better market efficiency. Here, each auction represents a CEX-DEX arbitrage opportunity. In the absense of revert protection, there is some chance auctions do not clear, leaving arbitrage opportunities unexploited and hence prices less accurate.
    • Revert protection offers better block space efficiency. Indeed, when there is revert protection, only a single, winning transaction consumes block space. On the other hand, without revert protection, all submitted transactions consume block space.
  • Our model allows different reversion penalty rates for the base fee and for priority fees. While the reversion penalty for priority fees influences bidder behavior, we establish that it does not influence aggregate system outcomes such as revenue, the probability that the auction clears, or the number of submitted transactions.
@@ 1.5 Literature Review
The dynamics of onchain MEV auctions were first explored in seminal work by Daian et al. [2]. Rasheed et al. mention the congestion effects of reverting transcations in [9]. Fox, Pai, and Resnick wrote about the feasibility of onchain auctions in [6], as did Robinson and White in [10]. The whitepaper for Unichain, an L2 that plans to offer revert protection, referred briefly to its efficiency benefits for decentralized exchanges [1].
Relative the economics literature on auctions, the model we consider is a common value, all pay auction, with a minimum allowable bid, and differential partial refunds for failed bids on the minimum bid and amounts above the minimum bid. Closest to our work in the paper of Hillman and Samet [7]. They consider a common value, all pay auction, with a minimum allowable bid, but do not allow for partial refunds of failed bids.
@@ 2.1 Auction Description
Each agent $i$ may choose to abstain or submit a bid (“priority gas fee”) $b_{i}\geq 0$; we denote the action of abstaining by $b_{i}=\varnothing$. The winner, denoted by $w$, extracts value $V$ from the MEV opportunity and pays $g+b_{w}$. In the event of a tie, the winner is randomly selected among the highest bids. Any losing agent $j$ does not receive any value and incurs a revert cost of $r_{1}g+r_{2}b_{j}$, where $r_{1},r_{2}\in[0,1]$ are the revert penalty rates on the base gas and priority gas fees, respectively. [^3] This model corresponds to a common value, all-pay auction with a minimum bid $g$ and differential refunds $(1-r_{1})g$ and $(1-r_{2})b_{j}$ for the base bid amount $g$ and additional bid amount $b_{j}$, respectively.
@@
Table 2: Revert Penalty Parameters for Various Settings
@@
In the cases involving MEV taxes in Table 2, observe that $r_{2}=0$. This is because with MEV taxes, applications can decide what fraction of the priority fees to capture themselves versus giving to the sequencer. Priority fees that are captured by the application are fully refunded on revert. As we will see later on (cf. Theorem 4.1), the total priority fee revenue does not depend on the choice of $r_{2}$. Hence, applications are incentivized to capture all but a negligible portion of the priority fee, and thus $r_{2}=0$.
We make the following assumption to avoid trivialities: [^4] This is because, if $V\leq g$, then the utility an agent receives when winning the action cannot be positive, even if the priority fee is zero, since the value does not exceed the base fee. Thus, all agents abstaining is a dominant strategy equilibrium.
@@ 2.2 Strategy Spaces and Payoffs
$\mathbb{P}(w=i|b)=\frac{\mathbbm{1}\{b_{i}=b_{M}\}}{\sum_{j\neq i}\mathbbm{1}\{b_{j}=b_{i}\}}.$
@@ 3.1 Pure Strategies
If $r_{1}=r_{2}=0$, and consider a pure strategy profile $b$. Assume the bids are ordered, so that $b_{1}\leq b_{2}\leq\dots\leq b_{N}$, with the possibility of abstaining represented by bids less than zero. Then $b$ is a Nash equilibrium if and only if $b_{N-1}=b_{N}=V-g$.
@@ 3.2 Mixed Strategies
for bids $b\in[0,V-g]$ and for each $i\in[N]$, and the expected payoff of every agent is zero.
@@
When $r_{1}>0$, the symmetric equilibrium strategy is characterized by a non-zero probability of abstention and a non-degenerate continuous CDF specifying the distribution from which to draw the priority gas fee bid. In the special case of $r_{1}=0$, all arbitrageurs participate with probability one. Figures 1 and 2 plot the equilibrium CDF and abstention probability for various configurations of parameters. In Appendix 0.B, we discuss comparative statics for the the equilibrium CDF and the abstention probability.
@@
We note that Theorem 3.3 is consistent with Theorem 3.1. When $r_{1}=r_{2}=0$, plugging these values into the formula in Theorem 3.3 yield $p^{*}=0$ and $F^{*}(b)=0$ for all $b\in[0,V-g)$, assuming that $g<V$, with $F^{*}(V-g)$ being undefined. Compare this with the pure strategy equilibrium given in Theorem 3.1, which corresponds to a symmetric mixed-strategy Nash equilibrium given by $p^{*}=0$, $F^{*}(b)=0$ for all $b\in[0,V-g)$, and $F^{*}(V-g)=1$. Indeed, one can show that $F^{*}(b)$ converges pointwise to the function $\mathbbm{1}\{b=V-g\}$ as $(r_{1},r_{2})\to(0,0)$.

### v2

@@ Abstract
Revert protection is a feature provided by some blockchain platforms that prevents users from incurring fees for failed transactions. We study the economic implications and benefits of revert protection in the context of priority gas auctions and maximal extractable value. We develop a model in which searchers bid for a top-of-block arbitrage opportunity under varying degrees of revert protection. This model applies to a broad range of settings, including bundle auctions on L1s and priority ordering sequencing rules on L2s. We quantify, in closed form, how revert protection improves equilibrium auction revenue, market efficiency, and blockspace efficiency.
@@ 1 Introduction and Background
Revert protection is a feature for blockchains that block builders and sequencers can provide where they exclude transactions that would otherwise fail, protecting users from paying fees for failed transactions. However, failed transactions would pay fees, which often accrue to the block builder or sequencer. So is it in their interest to offer revert protection?
In this paper, we argue that revert protection is indeed in their interest and quantify the value of revert protection on blockchain platforms, particularly on the auctions that many of those platforms use to capture and allocate maximal extractable value (MEV). We show that in various relevant settings, revert protection is beneficial for auctioneer revenue, as well as for other relevant outcomes like price discovery and blockspace efficiency.
The analysis is complicated by the wide range of rules — across the complex Ethereum block builder market and on a variety of L2s — for how transactions are selected for inclusion and ordered, how fees are collected, and who those fees go to, as well as how applications behave under those rules. We propose a model that can be parameterized to cover a wide variety of these settings, and use that model to solve for the equilibrium behavior of MEV searchers as a function of those parameters.
@@ 1.1 Revert Protection
On Ethereum and similar blockchains, users interact with smart contracts using atomic transactions. In the execution of a transaction, a smart contract may trigger a “revert” in the transaction, causing the entire transaction to fail. Normally, if the transaction is included, the user who sent the reverting transaction is still charged some transaction fee, even though the transaction has no other effect on the blockchain state. The fee charged is typically given by the product of gas price of and the gas used by the transaction before it either completes or reverts. The gas price can be separated into the “base fee”—a flat fee that must be paid by any transaction in the block—and the “priority fee”—an additional fee paid to the builder that often affects inclusion and ordering (particularly when blocks are full).
The gas used in the transaction is a function of how much of the transaction was executed—a transaction that reverts uses only some portion of the gas that it would have used had it succeeded. For example, on an automated market maker, typical reverting transactions may use only around 10–-20% of the gas that a successful transaction would use. “Revert protection” is thereby the feature that block builders and sequencers may choose to implement in which failed transactions are excluded entirely from blocks. This feature improves user experience, as users only have to pay fees if their transaction succeeds. However, it also has significant effects on the behavior of the profit-seeking bots known as “searchers,” which we now describe.
1.2 Block-Building Auctions
Public blockchains have abundant opportunities for MEV. As an example, one major type of MEV is the arbitraging of prices between decentralized and centralized exchanges, also called “CEX-DEX arbitrage,” as discussed in [4] and [13]. On these blockchains, independent actors known as “searchers” seek out MEV opportunities and compete to fill them.
Some block builders use different types of auctions to allocate these opportunities to searchers. For a given MEV opportunity, each transaction attempting to claim it can be thought of as a “bid.” Generally, only one transaction trying to claim a given opportunity will succed in extracting the value from it. Revert protection is therefore very relevant for these auctions, because it protects bidders on a given MEV opportunity from having to pay for failed bids if they do not win the opportunity. Two examples of block building auctions include the “bundle auctions” run by builders such as Flashbots on Ethereum mainnet and the “priority ordering” rule implemented by sequencers on Ethereum L2s.
@@ 1.2 Block-Building Auctions
On Ethereum mainnet, blocks are typically built by profit-maximizing “builders.” Many of these builders run “bundle auctions” in which they allow any searcher to submit transactions, and use those searchers’ bids as a factor when deciding which transactions to include and how to order them. [^1] Block builders also play the role of bidders themselves in the “MEV-Boost” auction [7], in which they bid to have their block included by the current proposer. As “bids” are complete blocks in PBS auctions, revert protection is not as relevant here, and we will mostly set this case aside for the purposes of this paper.
@@
Bids can be expressed in two ways for this auction: (i) through the priority fee on the transaction or (ii) through a direct payment as part of the logic of the transaction by calling coinbase.transfer [6]. The relevant difference between these methods of payment is that some portion of priority fees are paid even if the transaction reverts, whereas if the transfer is made as part of the transaction, then it will be conditional on whether the transaction succeeds.
@@
On Ethereum mainnet, the two components of gas price (base and priority fee) are paid in ETH and accrue to different users: base fees are burned, meaning they ultimately accrue to all ETH holders, rather than the builder; priority fees accrue to the block builder. [^2] These bids, whether paid via priority fee or Coinbase transfer, technically go to the proposer, not the builder who is assembling the block. However, since they reduce the amount that the builder has to pay to the proposer to win the MEV-Boost auction, we can think of these payments as a value transfer to the builder. Many builders, such as Flashbots [8], provide revert protection for transactions submitted to them, even though it is not required by the Ethereum protocol and they might receive more in transaction fees by including it. Our results in this paper help explain this choice by showing that it likely increases their expected revenue in equilibrium.
Priority Ordering.
@@
On Layer 2 blockchains today, blocks are typically built by a single sequencer, which often follows a deterministic algorithm for transaction inclusion and ordering. One of the most popular algorithms for this is “priority ordering,” in which transactions in a block are ordered in descending order of their gas price.
@@
Priority ordering can be thought of as an auction in which transactions bid with the discretionary part of their gas price (i.e. their priority fee) to be included earlier in the chain. By default, this means that certain kinds of MEV (such as top-of-block CEX-DEX arbitrage) will generally accrue to the sequencer through priority fees. This means that when transactions fail, they still pay some of their priority fee to the sequencer. However, there is a mechanism called MEV taxes [16] that applications can use to capture all but a negligible portion of the value that would otherwise be paid through priority fees. Since MEV taxes are paid as part of the transaction and revert if the transaction fails, these fees will only be paid if the transaction succeeds.
@@
Even within these two settings of L1 block builders and L2 priority-ordered sequencers, we now need to consider at least seven cases that differ in who fees go to and how much is paid when the transaction reverts. Table 1 shows the differences between these cases. Our model is general enough to capture the distribution of transaction fees and revert behavior for all of these cases. We can compute in closed-form the expected base fee and remaining fee (i.e. priority fee, Coinbase transfer, or MEV tax) for all listed settings as well. For the remainder of the paper, we will use the terms “builder” and “sequencer” interchangeably.
@@
Table 1: Transaction Fee Distribution and Revert Behavior for Various Settings.
1.3 Implementation Costs of Revert Protection
This paper is primarily concerned with the potential effects of revert protection, not with its implementation. We note that feasible revert protection is a difficult technical challenge: knowing whether a transaction will revert usually requires a builder to execute part of the transaction itself, which consumes computational resources. If the builder does not charge the sender anything for that reverted transaction, denial-of-service (DOS) attacks are possible by spamming the builder with an overwhelming number of reverting transactions.
There are some methods to mitigate this kind of attack. In some settings, the builder can use out-of-band spam prevention techniques (such as IP blocking) to make DOS attacks infeasible. For certain cases—including the auction-like use cases described above—it is also possible for the builder to determine statically whether a transaction will fail or not.
We abstract away from these challenges in our baseline model, assuming that the builder can implement revert protection with negligible cost. In our extensions, we study two schemes to deal with the implementation costs of revert protection where (i) costs are internalized by the sequencer and (ii) searchers pay a fixed amount for revert protection services.
@@ 1.4 Our Contributions
  • We introduce a novel and unified game-theoretic model that can be used to analyze revert protection in a variety of settings such as L1 block builders, L2 priority-ordered sequencers, MEV taxes, etc.
  • We characterize equilibria of our model in closed form, in terms of the model parameters: the value of the MEV opportunity, base fee, revert penalty parameters, and number of participating agents.
  • Using our model equilibrium, we can quantify the benefits of revert protection versus not offering revert protection:
    • Revert protection offers higher auction revenue. The auctioneer extracts the full value of the MEV opportunity when the auction clears. However, in the absence of revert protection, agents randomize when they participate, and there is a non-zero probability that the auction does not clear and value is lost. This results in reduced sequencer revenue.
    • In the context of automated market makers, revert protection offers better market efficiency. Here, the auction represents a CEX-DEX arbitrage opportunity. In the absense of revert protection, there is some chance that auctions do not clear, leaving arbitrage opportunities unexploited and hence prices less accurate.
    • Revert protection offers better blockspace efficiency. With revert protection, only a single, winning transaction consumes block space. On the other hand, without revert protection, all submitted transactions consume block space.
  • Our model allows for different revert penalty rates for the base fee and priority fees. While the penalty rate for priority fees influences bidder behavior, we show that it does not affect aggregate outcomes (e.g. revenue, number of submitted transactions).
  • When implementation costs are non-negligible but not too large, we show that charging a fixed fee to searchers for offering revert protection results in more auction revenue than builders internalizing the costs themselves.
@@ 1.5 Literature Review
We contribute to the literature on on-chain MEV auctions, first explored in the seminal work of Daian et al. [4]. Studies have focused on aspects in the L1 block building environment, such as censorship resistance [5], advantages for integrated searcher-builders [14], private mempools [3], and proposer-builder separation [2]. Other studies have focused on designing on-chain auctions and analyzing auction equilibria for L2 networks with various sequencing rules, including auctions where bidders are required to pay deposits [17], auctions under shared sequencing [12], taxes on MEV for applications [16], and the TimeBoost mechanism [11] proposed for Arbitrum [10] (a L2 network with a first-come first-serve sequencing rule). Relative to these studies, our paper offers a novel analysis on the effect of revert protection on on-chain auctions. Prior work has mentioned the congestion effects of reverting transactions in [15], and the whitepaper for Unichain, an L2 network that plans to offer revert protection, claims that it benefits swapping on decentralized exchanges [1].
In the context of the economic literature on auctions, our model is a common-value all-pay auction with a reserve price and differential partial refunds for failed bids on the minimum bid and amount above the minimum bid. Common-value auction models are frequently used in the analysis of on-chain auctions [2, 3, 12, 14] and in the traditional economics literature as well. Closest to our work is [9], which also considers a common-value all-pay auction with a reserve price, however does not allow for partial refunds on failed bids, and [18], whose model allows for a wide range of payoff and cost functions but has a more constrained strategy space for the bidders.
@@ 2.1 Auction Description
Agent $i$ may choose to abstain or submit a bid (“priority gas fee”) $b_{i}\geq 0$; we denote the action of abstaining by $b_{i}=\varnothing$. The winner, denoted by $w$, extracts value $V$ from the MEV opportunity and pays $g+b_{w}$. In the event of a tie, the winner is randomly selected among the highest bids. Any losing agent $j$ does not receive any value and incurs a revert cost of $r_{1}g+r_{2}b_{j}$, where $r_{1},r_{2}\in[0,1]$ are the revert penalty rates on the base gas and priority gas fees, respectively. [^3] This model corresponds to a common value, all-pay auction with a minimum bid $g$ and differential refunds $(1-r_{1})g$ and $(1-r_{2})b_{j}$ for the base bid amount $g$ and additional bid amount $b_{j}$, respectively.
@@
Table 2: Revert Penalty Parameters for Settings in Table 1
@@
In the cases involving MEV taxes in Table 2, observe that $r_{2}=0$. This is because with MEV taxes, applications can decide what fraction of the priority fees to capture themselves versus giving to the sequencer. Priority fees that are captured by the application are fully refunded on revert. As we will see later on (cf. Theorem 4.1), the total priority fee revenue does not depend on the choice of $r_{2}$. Hence, applications are incentivized to capture all but a negligible portion of the priority fee, and thus $r_{2}=0$. We make the following assumption to avoid trivialities: [^4] This is because, if $V\leq g$, then the utility an agent receives when winning the action cannot be positive, even if the priority fee is zero, since the value does not exceed the base fee. Thus, all agents abstaining is a dominant strategy equilibrium.
@@ 2.2 Strategy Spaces and Payoffs
$\mathbb{P}(w=i|b)=\frac{\mathbbm{1}\{b_{i}=b_{M}\}}{1+\sum_{j\neq i}\mathbbm{1}\{b_{j}=b_{i}\}}.$
@@ 3.1 Pure Strategies
If $r_{1}=r_{2}=0$, then a pure strategy profile $b$ where bids are ordered such that $b_{1}\leq b_{2}\leq\dots\leq b_{N}$ (with abstentions represented by bids less than zero) is a Nash equilibrium if and only if $b_{N-1}=b_{N}=V-g$.
@@ 3.2 Mixed Strategies
for bids $b\in[0,V-g]$ and agents $i\in[N]$. The expected payoff of every agent in equilibrium is zero.
@@
When $r_{1}>0$, the symmetric equilibrium strategy is characterized by a non-zero probability of abstention and a non-degenerate continuous CDF specifying the distribution from which to draw the priority gas fee bid. In the special case of $r_{1}=0$, all arbitrageurs participate with probability one. Figures 1 and 2 plot the equilibrium CDF and abstention probability for various parameters.  [^5] In Appendix 0.B, we discuss comparative statics for the the equilibrium CDF and the abstention probability.
@@
We note that Theorem 3.3 is consistent with Theorem 3.1. When $r_{1}=r_{2}=0$, Theorem 3.3 yields $p^{*}=0$ and $F^{*}(b)=0$ for all $b\in[0,V-g)$. Compare this with the pure strategy equilibrium given in Theorem 3.1, which corresponds to a symmetric mixed-strategy Nash equilibrium given by $p^{*}=0$, $F^{*}(b)=0$ for all $b\in[0,V-g)$, and $F^{*}(V-g)=1$. Indeed, one can show that $F^{*}(b)$ converges pointwise to the function $\mathbbm{1}\{b=V-g\}$ as $(r_{1},r_{2})\to(0,0)$.